An advertisement for a pineapple drink claims that the drinkcontains 15% pineapple juice. How much pure pineapple juice would have to be added to 8 quarts of the drink to obtain a mixture containing 50% pineapple juice?
Answer by kapilsinghi(68)

15% of 8 quants = pinaple juice
(100% - 15%)85% = others
so we have (15/100)*8 quants of pinaple juice available
and we have (85/100)*8 quants of others available
to make the mixture as 50% juice and 50% others
the quantity of juice should be equal to quantity of others
so quantity of juice and others in new mixture is (85/100)*8 each
so we need to add x quants of juice to the
present quantity to make it (85/100)*8
so (15/100)*8 + x = (85/100)*8
=> x = (85/100)*8 - (15/100)*8
=> x = (70/100)*8
=> x = (7/10)*8
=> x = (7*8)/10
=> x = 5.6 quants
5.6 quants of juice is to be added to the old mixture to make it 50-50

kapilsinghi123@gmail.com